OpenStudy (anonymous):

Given the quadratic function f(x) = x^2 – 14x + 49, find a value of x such that f(x) = 36

OpenStudy (anonymous):

x^2-14x-49 factors to (x-7)^2 so (x-7)^2=36 square root both sides x-7=6 x=13

OpenStudy (anonymous):

13,-1

OpenStudy (angela210793):

x^2-14x+49=36 x^2-14x+13=0 do u know how to solve it now?

OpenStudy (angela210793):

Damonky's solution's better :D

OpenStudy (anonymous):

13 and 1

OpenStudy (anonymous):

ignore the minus typo

OpenStudy (anonymous):

damonky solution is better ya but 1 is also the possible answer rite? substitute in 1 into the equation you will still get 36 .